Paulson Company issues 6%, four-year bonds, on December 31, 2015, with a par value of $200,000 and semiannual interest payments. Semiannual Period-End Unamortized Discount Carrying Value (0) 12/31/2015 $ 13,466 $ 186,534 (1) 6/30/2016 11,782 188,218 (2) 12/31/2016 10,098 189,902 Use the above straight-line bond amortization table and prepare journal entries for the following. (a) The issuance of bonds on December 31, 2015. (b) The first interest payment on June 30, 2016. (c) The second interest payment on December 31, 2016.
Answer #1
Concepts and reason

Bonds payable: Bonds payable is a type of instrument where by an amount is borrowed for a fixed period with fixed rate of interest. Bonds payable is shown as a liability in the books of issuer of bonds. Amount written on the face of bonds issued is known as par value. If the bonds are issued at a value less than the par, then the bonds are said to be issued at a discount. If the bonds are issued at a value more than the par, then the bonds are said to be issued at a premium.

Discount on bonds payable: The excess of par value over the carrying value of the bonds will be termed as discount on bonds payable. It is amortized over the period of issuance of bonds. Discount is amortized using two methods- Straight-line amortization method and effective interest amortization method.

Straight-line amortization: Straight-line amortization method is the method of amortizing discount on bonds payable. In this method, the discount on bonds payable is amortized by dividing the total discount with the term of the bonds.

Fundamentals

Transactions: Transactions are the events that happen in a business for a particular period. These transactions form base for the accounting. These transactions are recorded in the books of accounts which help to prepare the ledger accounts.

Journal entries: The transactions of an organization are recorded in the books of accounts through journal entries. Analyzing and journalizing of transactions is the second step in the accounting cycle. These journal entries are used to post the transactions into ledger.

Debit and credit rules: In a journal entry, accounts are debited and credited. In case of real accounts, journal entry is recorded using the principle “Debit what comes in and Credit what goes out”.

In case of personal accounts, journal entry is recorded using the principle “Debit the receiver and Credit the giver”.

In case of nominal accounts, journal entry is recorded using the principle “Debit all expenses and losses and Credit all revenues and gains”.

(a)

Record the journal entry for the issuance of bonds on December 31, 2015 as shown below:

Ref.
General Journal
Date
Account Titles and Explanation
Dec. 31, 2015 Cash
Discount on bonds payable
Bonds payable
(To recor

(b)

Record the journal entry for the first interest payment on June 30, 2016 as shown below:

Ref.
Credit
Debit
$7,684
General Journal
Date
Account Titles and Explanation
A
Jun. 30, 2016 Interest expense [$6,000 + $1,68

Note: Compute the cash value of semi-annual interest on bonds as shown below:

Cashpaid=Parvalueofbonds×Interest%×6months12months=$200,000×6%×6months12months=$6,000\begin{array}{c}\\{\rm{Cash paid = Par value of bonds }} \times {\rm{ Interest \% }} \times {\rm{ }}\frac{{6{\rm{ months}}}}{{12{\rm{ months}}}}\\\\ = {\rm{ \$ 200,000 }} \times {\rm{ 6\% }} \times {\rm{ }}\frac{{{\rm{6 months}}}}{{12{\rm{ months}}}}\\\\ = {\rm{ \$ 6,000}}\\\end{array}

(c)

Record the journal entry for the first interest payment on December 31, 2016 as shown below:

Credit
Ref. Debit
$7,684
General Journal
Date
Account Titles and Explanation
Dec 31, 2016 Interest expense [$6,000+ $1,684]
C

Note: Compute the cash value of semi-annual interest on bonds as shown below:

Cashpaid=Parvalueofbonds×Interest%×6months12months=$200,000×6%×6months12months=$6,000\begin{array}{c}\\{\rm{Cash paid = Par value of bonds }} \times {\rm{ Interest \% }} \times {\rm{ }}\frac{{6{\rm{ months}}}}{{12{\rm{ months}}}}\\\\ = {\rm{ \$ 200,000 }} \times {\rm{ 6\% }} \times {\rm{ }}\frac{{{\rm{6 months}}}}{{12{\rm{ months}}}}\\\\ = {\rm{ \$ 6,000}}\\\end{array}

Ans: Part a

Ref.
General Journal
Date
Account Titles and Explanation
Dec. 31, 2015 Cash
Discount on bonds payable
Bonds payable
(To recor

Part b

Ref.
Credit
Debit
$7,684
General Journal
Date
Account Titles and Explanation
A
Jun. 30, 2016 Interest expense [$6,000 + $1,68

Part c

Credit
Ref. Debit
$7,684
General Journal
Date
Account Titles and Explanation
Dec 31, 2016 Interest expense [$6,000+ $1,684]
C